9. Ordered: Lidocaine 1 mg/min IV Available: Lidocaine 2 gram/500 ml Calculate the number of ml/hr
Answer #1

9)Given value:- Ordered :- Lidocaine :- 1mg/min IV

Available :- Lidocaine 2 gram/500ml

1 gram = 1000 mg

2gram= 2000mg

2000mg Lidocaine contained in 500ml

1 mg Lidocaine = 500/2000 = 0.25 ml

In 1 minute, 0.25 ml of Lidocaine is intravenously infuse

So in 60 minutes, amount of Lidocaine intravenously infuse= 0.25 × 60 = 15 ml

Hence , 15 ml/ hr